A. Use the high-low method to answer the following: How much are variable times per unit? How much are total fixed times? Write the time equation in proper form. Then, draw a slope (m), and determine of the life cycle planning and life cycle integration Manufacturing MethodsUnits Times Operation1 Operation 2 Operation 3 Operation 4 Operation 5 1.208 1.113 1.183 1.203 1. 213 40/hours 39/ hours 3/ hours I/ hours 0/ hours
Answer #1

High Low method takes two extreme activity levels (i.e. labor hours, machine hours, etc.) from a set of actual data of various activity levels and their corresponding total cost figures.These figures are then used to calculate the approximate variable cost per unit (b) and total fixed cost (a) to obtain a cost volume formula:

y=a+bx

High-Low Method Formulas:-

A. Variable Cost per Unit

Variable cost per unit (b) is calculated using the following formula:

Variable Cost per Unit = y2 − y1
x2 − x1

Where,
y2 is the total unit at highest level of activity;
y1 is the total unit at lowest level of activity;
x2 are the number of units at highest level of activity; and
x1 are the number of units at lowest level of activity

The variable time per unit is equal to the slope of the cost volume line (i.e. change in total time ÷ change in number of units produced).

in given question it shall be as under:-

y2 =60/hrs

y1 =39/hrs

X2 =1,213 units

X1 =1,113 units

Variable time per Unit = (60 − 39) ÷ (1.213 − 1.113) = 210 hr per unit

Variable time per unit= 210 hr

B. Total Fixed Cost

Total fixed cost (a) is calculated by subtracting total variable cost from total cost, thus:

Total Fixed Cost = y2 − bx2 = y1 − bx1

here:-

Total Fixed time = 60 − (210*1.213) = 39 − (210 × 1.113) =

= -194.73

As fixed time is negative, which is not realistic we shall use second highest values now:-

y2 =40/hrs

y1 =39/hrs

X2 =1,208 units

X1 =1,113 units

Variable time per Unit = (40 − 39) ÷ (1.208 − 1.113) = 10.53 hr per unit

Total Fixed time = 40 − (10.53*1.208) = 39 − (10.53× 1.113) = 27.28 hrs

Slope of "m"

Y=mx+b

where Y = total cost

          m = the slope of the line, i.e., unit variable cost

          X = the number of units of activity

          b = the y-intercept, i.e., total fixed costs

here:- unit variable time = 10.53 hr per unit